# references.bib

### v1

@@ 5 Ren–Xu counterexamples
for arbitrary $f_{i}$ and $h_{j}$. We know for any choice of $f_{i}$ and $h_{j},$ we have $P,Q\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$ by Lemma 5.14.
We claim that $B|\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]/(x_{1}-x_{2})^{2}.$ By [wang2022explicit] $A^{\prime}$ and $B^{\prime}$ are both polynomials in the variables $(x_{1}-x_{2})^{2}$ and $(x_{1}-x_{3})(x_{2}-x_{3})$. Moreover, by Lemma 5.9, $(x_{1}-x_{2})^{2}\nmid B$ so $B\equiv cM_{3m+1}\pmod{(x_{1}-x_{2})^{2}}$ for some $c\in\mathbf{F}_{3}$ such that $c\neq 0$. Similarly we know $\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\equiv c^{\prime}M_{3m+1}\pmod{(x_{1}-x_{2})^{2}}$ for some $c^{\prime}\in\mathbf{F}_{3}$. Thus we have $\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)=dB$ where $d=\frac{c^{\prime}}{c}.$
Note that
$(x_{1}+x_{2}+x_{3})M_{2i}$ $=(x_{1}+x_{2}+x_{3})(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=M_{2i+1}$
and
$(x_{1}+x_{2}+x_{3})M_{2i+1}$ $=(x_{1}+x_{2}+x_{3})^{2}(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=((x_{1}-x_{2})^{2}+(x_{1}-x_{3})(x_{2}-x_{3}))(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=(x_{1}-x_{3})(x_{2}-x_{3})M_{2i}$
$=M_{2i+2}.$
Using this we have
$PC+QB$ $=\left(h_{0}M_{5}B+f_{0}(x_{1}+x_{2}+x_{3})M_{4}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)x_{3}^{0}\right)$
$\quad+\sum_{i=1}^{3}\left(h_{i}M_{5-i}Bx_{3}^{i}+f_{i}(x_{1}+x_{2}+x_{3})M_{4-i}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)x_{3}^{i}-f_{i-1}M_{5-i}Bx_{3}^{i}\right)$
$\quad+h_{4}M_{1}Bx_{3}^{4}-f_{3}M_{1}Bx_{3}^{4}$
$=\left(h_{0}B+f_{0}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\right)M_{5}$
$\quad+\sum_{i=1}^{3}\left(\left(\left(h_{i}-f_{i-1}\right)B+f_{i}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\right)M_{5-i}x_{3}^{i}\right)+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
$=\left(h_{0}+f_{0}d\right)BM_{5}+\sum_{i=1}^{3}\left(\left(h_{i}-f_{i-1}\right)+f_{i}d\right)BM_{5-i}x_{3}^{i}+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
Letting $h_{i}$ be arbitrary for $i>0$, set $f_{3}=h_{4}$, $f_{i-1}=h_{i}+f_{i}d$ for $30<i<$ 3 and set $h_{0}=-f_{0}d$. This makes the expression $PC+QB=0.$
@@
If $(PC+QB)(x_{1}-x_{2})^{2m+1}$ were generated by $R_{m+1},$ there exists a symmetric polynomial $I$ such that $IR_{m+1}=(PC+QB)(x_{1}-x_{2})^{2m+1}.$ This implies $(I(x_{1}+x_{2}+x_{3})-P)C+(IS-Q)B=0.$ If $I(x_{1}+x_{2}+x_{3})-P\neq 0$ or $IS-Q\neq 0,$ there is a relation on $C$ and $B$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$, but $C(x_{1}-x_{2})^{2m+1}$ and $B(x_{1}-x_{2})^{2m+1}$ are generators of $Q_{m}(3,\mathbf{F}_{3}).$ Thus we must have $P=I(x_{1}+x_{2}+x_{3}),$ so $(x_{1}+x_{2}+x_{3})|P.$ Now we consider the symmetric polynomials $P^{\prime}=P+e_{2}^{2}+e_{2}e_{1}^{2}+e_{1}^{4}$ and $Q^{\prime}=Q+e_{3}e_{1}^{2}+(-d-1)e_{2}^{2}e_{1}-de_{1}^{3}e_{2}+(-d+1)e_{1}^{5}$. In $F_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2},$ we get that
@@
Then $f_{2}+1=(h_{3}+f_{3}d)+1=(h_{3}+1)+f_{3}d,$ $f_{1}=h_{2}+f_{2}d=(h_{2}-d)+(f_{2}+1)d,$ and the rest of the equations necessary for $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}$ are the same as $PC+QB\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Thus $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Moreover, $(x_{1}+x_{2}+x_{3})$ divides into $P+e_{2}e_{1}^{2}+e_{1}^{4}$ but not $e_{2}^{2},$ so $(x_{1}+x_{2}+x_{3})\nmid P^{\prime}.$ We have shown that if $(PC+QB)(x_{1}-x_{2})^{2m+1}$ is generated by $L,$ then $(x_{1}+x_{2}+x_{3})|P,$ implying $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is not generated by $R_{m+1}$. If $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is not a generator, then whatever generates it violates Lemma 4.3, so $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is indeed a degree $63m+$ 6 generator of $Q_{m+1}(3,\mathbf{F}_{3}).$
@@
Let $m\not\in X$ for some natural number $m.$ Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m+1}(3,\mathbf{F}_{3})$ and $L$ is another generator in degree $63m+$ 6 . Further let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$ where $d$ is maximal. Then $L\prod(x_{i}-x_{j})^{2(i-1)},$ $R_{m+1},$ and $1$ freely generate $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ for $1\leq i\leq d.$
@@
$L\prod(x_{i}-x_{j})^{2(i-1)}$ lies in a copy of $\mathrm{sign-triv}$ and it is divisible by $(x_{1}-x_{2})^{2(m+i)+1},$ so it must lie in $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $3m+6i$ by Lemma 4.1. If $L\prod(x_{i}-x_{j})^{2(i-1)}$ is not a generator, $R_{m+1}$ must generate $L\prod(x_{i}-x_{j})^{2(i-1)},$ implying a relation between $R_{m+1}$ and $L.$ Thus $L\prod(x_{i}-x_{j})^{2(i-1)}$ is indeed a generator.
Moreover, $33m+3+3m+6i=6(m+i)+$ 3 so by Lemma 4.4, $L\prod(x_{i}-x_{j})^{2(i-1)}$ and $R_{m+1}$ generate $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
Next, we prove that, for all consecutive spaces of quasi-invariants in the second half of the staircase, the lower degree generator is $\prod(x_{i}-x_{j})^{2}$ times the previous lower degree generator.
@@
Let $m\not\in X$ for some natural number $m.$ Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m}(3,\mathbf{F}_{3})$ and $L$ is another generator in degree $63m+$ 6 . Let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$ where $d$ is maximal. Further, let $L$ have degree $5$ in $x_{3}$. Then for all $d\leq i<2d,$ $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $3m+6d,$ $R_{m+1}\prod(x_{i}-x_{j})^{2(i-d)}$ in degree $3m+6(i-d)+3,$ and $1$
@@
The generator $R_{m+1}$ of $Q_{m+d}(3,\mathbf{F}_{3})$ is in degree $3m+3=3m+6(d-d)+3,$ and from Lemma 5.16 a second generator is $L\prod(x_{i}-x_{j})^{2(d-1)}$ in degree $3m+6d.$ Moreover these are the only generators so the claim is true for $i=d.$
Let $j$ be a natural number with $d<j<2d$ and suppose $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m+6d$ and degree $33m+6(i-d)+$ 3 for all $d\leq i<j$ where this upper degree generator is a polynomial of degree at most $5$ in $x_{3}$ and is not generated by $R_{m+1}$. Consider $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$ We know $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is an element of $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $33m+6(j-d-1)+$ 3 by Lemma 4.1. Since $j-1<j$ we may use our inductive hypothesis implies $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is a generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$.
Let $T$ be the degree $3m+6d$ generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $5$ in $x_{3}.$ We write $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}=R_{m+1}^{\prime}(x_{1}-x_{2})^{2(m+j-1)+1}$ and $T=T^{\prime}(x_{1}-x_{2})^{2(m+j-1)+1}$ for $s_{12}$ invariant polynomials $R_{m+1}^{\prime}$ and $T^{\prime}$. If $o=m+4j-6d-2$ and $r=m+6d-2j+1,$ then $\deg R_{m+1}^{\prime}=o$ and $\deg T^{\prime}=r.$ We want to find a degree $r-o$ symmetric polynomial $P$ such that
@@
We claim that $R_{m+1}^{\prime}$ is degree 00 in $x_{3}.$ This is because $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}=P_{l}^{3^{a}}\prod(x_{i}-x_{j})^{2(j-d-1)}$ as we proved in Lemma 5.12. Since $P_{l}$ is the map of the generator of $Q_{l}(3,\mathbf{Q})$ into characteristic $3,$ $P_{l}$ must be constant in the variable $x_{3}.$ We can see $\prod(x_{i}-x_{j})^{2(j-d-1)}$ is also constant in $x_{3},$ so $R_{m+1}$ and $R_{m+1}^{\prime}$ are constant in $x_{3}.$

### v2

@@ 1. In $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, for $j\in Z_{\geq 0}$,
for arbitrary $f_{j}$ and $h_{j}$. By Lemma 5.15, we know that for any choice of $f_{j}$ and $h_{j}$, we have $P,Q\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$.
We claim that $B|\pi\big{(}\frac{A^{\prime}-B^{\prime}}{3}\big{)}$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]/(x_{1}-x_{2})^{2}$. By [13], $A^{\prime}$ and $B^{\prime}$ are both polynomials in the variables $(x_{1}-x_{2})^{2}$ and $(x_{1}-x_{3})(x_{2}-x_{3})$. Moreover, by Lemma 5.9, $(x_{1}-x_{2})^{2}\nmid B$ so $B\equiv cM_{m}\pmod{(x_{1}-x_{2})^{2}}$ for some $c\in\mathbf{F}_{3}$ such that $c\neq 0$. Similarly, we know $\smash{\pi\big{(}\frac{A^{\prime}-B^{\prime}}{3}\big{)}}\equiv c^{\prime}M_{m}\pmod{(x_{1}-x_{2})^{2}}$ for some $c^{\prime}\in\mathbf{F}_{3}$. Thus we have $\smash{\pi\big{(}\frac{A^{\prime}-B^{\prime}}{3}\big{)}}=dB$, where $d=\frac{c^{\prime}}{c}$.
We use Lemma 5.13 to expand $PC+BQ$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]/(x_{1}-x_{2})^{2}$,
$PC+QB$ $=\biggl{(}h_{0}M_{5}B+f_{0}(x_{1}+x_{2}+x_{3})M_{4}\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}x_{3}^{0}\biggr{)}$
$\quad+\sum_{j=1}^{3}\biggl{(}h_{j}M_{5-j}Bx_{3}^{j}+f_{j}(x_{1}+x_{2}+x_{3})M_{4-j}\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}x_{3}^{j}$
${}-f_{j-1}M_{5-j}Bx_{3}^{j}\bigg{)}+h_{4}M_{1}Bx_{3}^{4}-f_{3}M_{1}Bx_{3}^{4}$
$=\biggl{(}h_{0}B+f_{0}\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}\biggr{)}M_{5}$
$\quad+\sum_{j=1}^{3}\biggl{(}\biggl{(}(h_{j}-f_{j-1})B+f_{j}\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}\biggr{)}M_{5-j}x_{3}^{j}\biggr{)}+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
$=(h_{0}+f_{0}d)BM_{5}+\sum_{j=1}^{3}\biggl{(}(h_{j}-f_{j-1})+f_{j}d\biggr{)}BM_{5-j}x_{3}^{j}+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
Letting $h_{j}$ be arbitrary for $j>0$, set $f_{3}=h_{4}$, $f_{j-1}=h_{j}+f_{j}d$ for $30<j<$ 3 and set $h_{0}=-f_{0}d$. This makes the expression $PC+QB=0$.
@@
If $(PC+QB)(x_{1}-x_{2})^{2m+1}$ were generated by $R_{m+1}$, there would exist a symmetric polynomial $I$ such that $IR_{m+1}=(PC+QB)(x_{1}-x_{2})^{2m+1}$. This implies $(I(x_{1}+x_{2}+x_{3})-P)C+(IS-Q)B=0$. If $I(x_{1}+x_{2}+x_{3})-P\neq 0$ or $IS-Q\neq 0$, there is a relation on $C$ and $B$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$, but $C(x_{1}-x_{2})^{2m+1}$ and $B(x_{1}-x_{2})^{2m+1}$ are generators of $Q_{m}(3,\mathbf{F}_{3})$. Thus we must have $P=I(x_{1}+x_{2}+x_{3})$, so $(x_{1}+x_{2}+x_{3})|P$. Now we consider the symmetric polynomials $P^{\prime}=P+e_{2}^{2}+e_{2}e_{1}^{2}+e_{1}^{4}$ and $Q^{\prime}=Q+e_{3}e_{1}^{2}+(-d-1)e_{2}^{2}e_{1}-de_{1}^{3}e_{2}+(-d+1)e_{1}^{5}$. In $F_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, we get that
@@
Then $f_{2}+1=(h_{3}+f_{3}d)+1=(h_{3}+1)+f_{3}d$, $f_{1}=h_{2}+f_{2}d=(h_{2}-d)+(f_{2}+1)d$, and the rest of the equations necessary for $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}$ are the same as $PC+QB\equiv 0\pmod{(x_{1}-x_{2})^{2}}$. Thus $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}$. Moreover, $(x_{1}+x_{2}+x_{3})$ divides into $P+e_{2}e_{1}^{2}+e_{1}^{4}$ but not $e_{2}^{2}$, so $(x_{1}+x_{2}+x_{3})\nmid P^{\prime}$. We have shown that if ${(PC+QB)(x_{1}-x_{2})^{2m+1}}$ is generated by $R_{m+1}$, then $(x_{1}+x_{2}+x_{3})|P$, implying ${(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}}$ is not generated by $R_{m+1}$. If $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is not a generator, then whatever generates it violates Lemma 4.3, so $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is indeed a degree $63m+$ 6 generator of $Q_{m+1}(3,\mathbf{F}_{3})$. ∎
@@
Let $m\not\in X$ for some natural number $m$. Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ and $L$ is another generator in degree $63m+$ 6. Further, let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$, where $d$ is maximal. Then $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$, $R_{m+1}$, and $1$ freely generate $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ for $1\leq j\leq d$.
@@
As a generator, $L$ lies in a copy of $\mathrm{sign-triv}$ and is divisible by $(x_{1}-x_{2})^{2(m+1)+1}$ by Lemma 4.1. Since $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$ is divisible by $(x_{1}-x_{2})^{2(m+j)+1}$, by the second part of Lemma 4.1, $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$ is in $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. If $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$ is not a generator, $R_{m+1}$ must generate $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$, implying a relation between $R_{m+1}$ and $L$. Thus $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$ is indeed a generator.
Moreover, $33m+3+3m+6j=6(m+j)+$ 3 so by Lemma 4.4, $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-1)}$ and $R_{m+1}$ generate $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. ∎
Next, we prove that, for all consecutive spaces of quasi-invariants in the second half of the staircase, the lower degree generator is $\prod(x_{i_{1}}-x_{i_{2}})^{2}$ times the previous lower degree generator.
@@
Let $m\not\in X$ for some natural number $m$. Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m}(3,\mathbf{F}_{3})$ and $L$ is another generator in degree $63m+$ 6. Let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$, where $d$ is maximal. Further, let $L$ have degree at most $5$ in $x_{3}$. Then for all $d\leq j<2d$, $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $3m+6d$, $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(j-d)}$ in degree $33m+6(j-d)+$ 3, and $1$.
@@
The generator $R_{m+1}$ of $Q_{m+d}(3,\mathbf{F}_{3})$ is in degree $33m+3=3m+6(d-d)+$ 3, and from Lemma 5.17 a second generator is $L\prod(x_{i_{1}}-x_{i_{2}})^{2(d-1)}$ in degree $3m+6d$. Moreover, these are the only generators so the claim is true for $j=d$.
Let $k$ be a natural number with $d<k<2d$ and suppose $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m+6d$ and degree $33m+6(j-d)+$ 3 for all $d\leq j<k$, where this upper degree generator is a polynomial of degree at most $5$ in $x_{3}$ and is not generated by $R_{m+1}$. Consider $Q_{m+k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. We know $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$ is an element of $Q_{m+k-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $33m+6(k-d-1)+$ 3 by Lemma 4.1. Since $k-1<k$, our inductive hypothesis implies $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$ is a generator for $Q_{m+k-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$.
Let $T$ be the degree $3m+6d$ generator for $Q_{m+k-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $5$ in $x_{3}$. We write $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}=R_{m+1}^{\prime}(x_{1}-x_{2})^{2(m+k-1)+1}$ and $T=T^{\prime}(x_{1}-x_{2})^{2(m+k-1)+1}$ for $s_{12}$ invariant polynomials $R_{m+1}^{\prime}$ and $T^{\prime}$. If $o=m+4k-6d-2$ and $r=m+6d-2k+1$, then $\deg R_{m+1}^{\prime}=o$ and $\deg T^{\prime}=r$. We want to find a degree $r-o$ symmetric polynomial $P$ such that
@@
We claim that $R_{m+1}^{\prime}$ has degree 00 in $x_{3}$. This is because $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}=P_{l}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$ as we proved in Lemma 5.12. Since $P_{l}$ is the map of the generator of $Q_{l}(3,\mathbf{Q})$ into characteristic $3$, $P_{l}$ must be constant in the variable $x_{3}$. We can see ${\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}}$ is also constant in $x_{3}$, so $R_{m+1}$ and $R_{m+1}^{\prime}$ are constant in $x_{3}$.